OpenStudy (anonymous):

a recipe that makes 6 servings calls for 1/3 cup of orange juice and 1/2 cup of lemonade. arlene modifies the recipe so that it can serve 9.

OpenStudy (carolina4567):

1/3 cup -> 6 servings ? cup -> 9 servings Rule of three : 9 * (1/3) / 6 = 1/2 cup

OpenStudy (igreen):

Well first you have to find out what half of 1/3 is, then add it to 1/3. Then you have to find out what half of 1/2 is, then add it to 1/2.

OpenStudy (carolina4567):

1/2 Cups

OpenStudy (igreen):

Half of 1/3 is 1/6. And half of 1/2 is 1/4. 1/3 + 1/6 = 1/2 1/2 + 1/4 = 3/4 So Arlene needs 1/2 cup of orange juice and 3/4 cup of lemonade.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @allymae2002

OpenStudy (texaschic101):

6/(1/3) = 9/x -- (6 serv to 1/3 cup oj = 9 serv to x cup oj) cross multiply (6)(x) = (1/3)(9) 6x = 3 x = 1/2 she would need 1/2 cups of oj to make 9 servings 6/(1/2) = 9/x --(6 serv to 1/2 cup lem = 9 serv to x cups lem) cross multiply (6)(x) = (1/2)(9) 6x = 9/2 x = (9/2) / 6 x = 9/2 * 1/6 x = 9/12 which reduces to 3/4 she would need 3/4 cup of lemonade to make 9 servings